OpenStudy (anonymous):

log5 4+log5 (2x)=log5 24

Nnesha (nnesha):

familiar with the log properties ?

OpenStudy (anonymous):

nope

Nnesha (nnesha):

have a look!quotient rule\[\large\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (anonymous):

but whats the answer!!

Nnesha (nnesha):

what property you should apply first ?

OpenStudy (netlopes1):

\[\log_5 {4} + \log_5 {2x} = \log_5 {24}\]\[\log_5 {8x} = \log_5 {24}\rightarrow 8x=24 \rightarrow x=3\]

OpenStudy (netlopes1):

Remember that \[x>0 \rightarrow x=3 \quad is \quad the \quad solution\]